# Reality conditions for the KdV equation and quasi-periodic solutions in finite phase spaces

### v2

@@ 5.3. Summary
The uniformization of the spectral curve is given by (5.1) in an implicit form. On the other hand, it brings explicit expressions (5.1) for dynamic variables $\beta_{2n-1}$, $\alpha_{2n}$. Coordinates $u_{1}$ and $u_{3}$ of the Jacobian variety of the spectral curve serve, up to a constant multiple $-\mathrm{b}$, as parameters $x$ and $t$ of the stationary and evolutionary flows, correspondingly.
@@ 6.1. Singularities of Weierstrass-$\wp$-functions
(62) $\begin{split}&\{2k-1\}\sim\tfrac{1}{2}\omega^{\prime}_{k}+\sum_{i=1}^{k-1}\tfrac{1}{2}\omega_{i},\quad\{2k\}\sim\tfrac{1}{2}\omega^{\prime}_{k}+\sum_{i=1}^{k}\tfrac{1}{2}\omega_{i},\quad k=1,\dots,g,\\ &\{2g+1\}=\sum_{k=1}^{g}\tfrac{1}{2}\omega_{k}.\end{split}$
Every half-period has the form $\Omega(I)+\Omega^{\prime}(I^{\prime})$, where $\Omega(I)$ is generated from $\tfrac{1}{2}\omega_{k}$, $k=1$, …, $g$, and $\Omega^{\prime}(I^{\prime})$ is generated from $\tfrac{1}{2}\omega^{\prime}_{k}$, $k=1$, …, $g$. Namely,
@@
And the sigma function does not vanish at these half-periods. Within the fundamental domain, the only zero of the sigma function is located at 00.
If $g>1$, then each $\mathcal{I}_{0}$ is obtained from $g$ sets $\{\iota\}$ with all indices different. The corresponding $I$ is obtained by taking the union of $g$ subsets from the collection $\mathfrak{I}_{1}$, and dropping indices which occur even number of times. If at least two subsets in this union coincide, then the resulting $I^{\prime}$ can be obtained as a union of $g-2$ or less number of subsets. That means, that the subspace $\mathfrak{U}_{\Re}(I^{\prime})$ contains zeros of the sigma function. Thus, the required $\hat{I}^{\prime}$ is obtained by the union of $g$ different and not empty subsets from the collection $\mathfrak{I}^{\prime}_{1}$. This implies $\hat{I}^{\prime}=\{1,2,\dots,g\}$. ∎
@@
$\Omega(\hat{I})=\left\{\begin{array}[]{ll}\sum_{k=1}^{\mathfrak{k}}\tfrac{1}{2}\omega_{2k},&g=2\mathfrak{k}\\ \sum_{k=0}^{\mathfrak{k}}\tfrac{1}{2}\omega_{2k+1}&g=2\mathfrak{k}+1,\end{array}\right.$
@@ 6.2. Hyperelliptic addition law
Let $\Upsilon_{n}(u)$, $n=2$, …, $g+2$, be $g$-component vector-functions of $u\in\operatorname{\mathfrak{J}}$. We introduce a matrix-function $\textsf{Q}(u)=\big{(}\Upsilon_{2}(u)$, $\Upsilon_{3}(u)$, …, $\Upsilon_{g+1}(u)\big{)}$ with entries $(\textsf{q}_{i,j}(u))$, and a vector-function $\bm{q}(u)=\Upsilon_{g+2}(u)$. Actually,
$\Upsilon_{2}(u)\equiv\big{(}\textsf{q}_{i,1}(u)\big{)}=\big{(}\wp_{1,2i-1}(u)\big{)},\quad i=1,\dots,g,$
$\Upsilon_{3}(u)\equiv\big{(}\textsf{q}_{i,2}(u)\big{)}=-\tfrac{1}{2}\big{(}\wp_{1,1,2i-1}(u)\big{)},$
$\Upsilon_{k+1}(u)\equiv(\textsf{q}_{i,k}(u))=\textsf{q}_{1,k-2}(u)\Upsilon_{2}(u)+\Upsilon^{\circ}_{k-1}(u),\quad k=3,\dots,g+2,$
where $\text{where }\Upsilon^{\circ}_{k-1}(u)=\big{(}\textsf{q}_{2,k-2}(u),\textsf{q}_{3,k-2}(u),\dots,\textsf{q}_{g,k-2}(u),0\big{)}^{t}.$
Let $\bar{\eta}=(\eta_{3g},\dots,\eta_{g+4},\eta_{g+2})^{t}$, and $\eta=(\eta_{g},\dots,\eta_{2},\eta_{1})^{t}$. Let $u$, $v$, $w\in\operatorname{\mathfrak{J}}$, provided $u+v+w=0$. Then the system
$\begin{pmatrix}1_{g}&\textsf{Q}(u)\\ 1_{g}&\textsf{Q}(v)\\ 1_{g}&\textsf{Q}(w)\end{pmatrix}\begin{pmatrix}\bar{\eta}\\ \eta\end{pmatrix}=\begin{pmatrix}\bm{q}(u)\\ \bm{q}(v)\\ \bm{q}(w)\end{pmatrix}$
@@
$\textsf{P}=\big{(}\Upsilon_{2}(u)+\Upsilon_{2}(v),\Upsilon_{4}(u)+\Upsilon_{4}(v),\dots,\Upsilon_{2g}(u)+\Upsilon_{2g}(v)\big{)},$
where $\Upsilon_{k}$, $k>g+2$, are computed by the same rule
$\Upsilon_{k+1}(u)\equiv(\textsf{q}_{i,k}(u))=\textsf{q}_{1,k-2}(u)\Upsilon_{2}(u)+\Upsilon^{\circ}_{k-1}(u).$
Let a vector $\Pi=(\Pi_{0}$, $\Pi_{1}$, …, $\Pi_{g})$ be defined as follows
@@
and a vector $\textsf{H}=(\textsf{H}_{0}$, $\textsf{H}_{1}$, …, $\textsf{H}_{g})$, depending on the parity of $g$, be
(65b) $\begin{split}&g=2\mathfrak{k}-1:\qquad\textsf{H}_{k}=\sum_{i=1}^{k}\eta_{2k-2i+1}\eta_{2i-1}-\sum_{j=0}^{k}\lambda_{2k-2j}\sum_{i=0}^{j}\eta_{2j-2i}\eta_{2i},\\ &g=2\mathfrak{k}:\qquad\quad\ \ \textsf{H}_{k}=-\sum_{i=0}^{k}\eta_{2k-2i}\eta_{2i}+\sum_{j=1}^{k}\lambda_{2k-2j}\sum_{i=1}^{j}\eta_{2j-2i+1}\eta_{2i-1},\end{split}$
provided $\eta_{0}=1$, and $\lambda_{0}=1$. Then addition formulas for $\wp_{1,2i-1}$ are given by
(66) $\wp_{1,2i-1}(w)=\sum_{j=0}^{i}\textsf{H}_{i-j}\Pi_{j}.$
They work in an arbitrary genus $g$, and $\eta_{k}$ is supposed to be zero if there is no such entry in $\bar{\eta}$, $\eta$ in this genus. In particular,
(67) $\begin{split}&g=2\mathfrak{k}-1:\qquad\wp_{1,1}(w)=\eta_{1}^{2}-2\eta_{2}-\lambda_{2}-\wp_{1,1}(u)-\wp_{1,1}(v),\\ &g=2\mathfrak{k}:\qquad\quad\ \ \wp_{1,1}(w)=\eta_{1}^{2}-2\eta_{2}-\wp_{1,1}(u)-\wp_{1,1}(v),\end{split}$
where $\eta_{2}=0$ in the case of genus $1$.
Expressions for $\eta_{k}$ in terms of $\wp_{1,2i-1}(u)$, $\wp_{1,2i-1}(v)$, $\wp_{1,1,2i-1}(u)$, $\wp_{1,1,2i-1}(v)$ are obtained from the system
(74) $\begin{pmatrix}1_{g}&\textsf{Q}(u)\\ 1_{g}&\textsf{Q}(v)\end{pmatrix}\begin{pmatrix}\bar{\eta}\\ \eta\end{pmatrix}=\begin{pmatrix}\bm{q}(u)\\ \bm{q}(v)\end{pmatrix}.$
Finally, expressions for $\wp_{1,1,2i-1}(w)$ in terms of $\wp_{1,2i-1}(w)$ and $\eta_{k}$ are obtained from
(75) $\bar{\eta}+\textsf{Q}(w)\eta=\bm{q}(w),$
@@ 6.3. Real-valued Weierstrass-$\wp$-functions
Through the help of the addition law one can obtain expressions for $\wp_{1,2i-1}(u+\imath\upsilon)$ and $\wp_{1,1,2i-1}(u+\imath\upsilon)$ in terms of $\wp_{1,2i-1}(u)$, $\wp_{1,2i-1}(\imath\upsilon)$, $\wp_{1,1,2i-1}(u)$, $\wp_{1,1,2i-1}(\imath\upsilon)$, $i=1$, …, $g$. The mentioned $2g$ Weierstrass-$\wp$-functions serve as generators in the differential field of all multiply periodic functions on a hyperelliptic curve of genus $g$, see [10]. These $2g$ functions arise in the solution (2.3) of the Jacobi inversion problem.
Next, we prove that $\wp_{i,j}$ on subspaces $\mathfrak{U}_{\Re}(I^{\prime})$, and $\mathfrak{U}_{\Im}(I)$ are real-valued.
@@
Suppose $\wp_{1,1,2i-1}(\imath\upsilon)=0$. Then for all $u\in\operatorname{\mathbb{R}}^{g}$ functions $\wp_{1,2i-1}(u+\imath\upsilon)$ and $\wp_{1,1,2i-1}(u+\imath\upsilon)$, $i=1$, …, $g$, are real-valued.
@@
Recall, that $\wp_{i,j}$ are even functions, and so $\wp_{i,j}(\imath\upsilon)=\widehat{\wp}_{i,j}(\upsilon)$ are real-valued. On the other hand, $\wp_{i,j,k}$ are odd, and so $\wp_{i,j,k}(\imath\upsilon)=\imath\widehat{\wp}_{i,j,k}(\upsilon)$ are purely imaginary. Since $\imath$ arises only from $\wp_{1,1,2i-1}(\imath v)$, and $\wp_{1,1,2i-1}(\imath v)=0$, all functions $\wp_{1,2i-1}(u+\imath v)$ and $\wp_{1,1,2i-1}(u+\imath v)$ are real-valued. ∎
@@
Suppose $\wp_{1,1,2i-1}(u)=0$. Then for all $\upsilon\in\operatorname{\mathbb{R}}^{g}$ functions $\wp_{1,2i-1}(u+\imath\upsilon)$, $i=1$, …, $g$, are real-valued, and $\wp_{1,1,2i-1}(u+\imath\upsilon)$, $i=1$, …, $g$, are purely imaginary.
@@
From the structure of Q and $\bm{q}$ in (74), we see the following.
@@
If $g=2\mathfrak{k}-1$, then Q contains $\mathfrak{k}-1$ columns linear in odd functions $\wp_{1,1,2i-1}$, and $\bm{q}$ is linear in $\wp_{1,1,2i-1}$. These columns of $\textsf{Q}(\imath v)$ and $\bm{q}(\imath v)$ have purely imaginary entries. Due to $\wp_{1,1,2i-1}(u)=0$, the corresponding columns of $\textsf{Q}(u)$ vanish, as well as $\bm{q}(u)$. The system (74) is solved by Cramer’s rule. Thus, all entries of $\bar{\eta}$, which are indexed by odd numbers, have purely imaginary values, as well as entries $\eta_{k}$ of $\eta$ with odd $k$. Entries $\eta_{k}$ with even $k$ are real-valued.
@@
If $g=2\mathfrak{k}$, then Q contains $\mathfrak{k}$ columns linear in odd functions $\wp_{1,1,2i-1}$, and $\bm{q}$ is expressed in terms of even functions $\wp_{1,2i-1}$ only. Thus, all entries of $\bar{\eta}$, which are indexed by even numbers, are real-valued, as well as entries $\eta_{k}$ of $\eta$ with even $k$. Entries $\eta_{k}$ with odd $k$ are purely imaginary. Therefore, $\eta_{k}$ with odd $k$ are purely imaginary, and with even $k$ are real-valued, in any genus.
@@
Next, we look at the expressions for $\textsf{H}_{k}$ given by (65b). Evidently, all $\textsf{H}_{k}$ are real-valued, since $\eta_{k}$ with odd $k$ appear in terms $\eta_{2j-2i+1}\eta_{2i-1}$, which are real-valued. Therefore, expressions for $\wp_{1,2i-1}(u+\imath v)$ given by (66) are real-valued.
@@
Finally, we look at (75). If $g=2\mathfrak{k}-1$, then $\eta_{k}$ with odd $k$ are multiplied by real-valued columns expressed in terms of even functions $\wp_{1,2i-1}(w)$ only. Since $\eta_{k}$ with odd $k$ are purely imaginary, the expression $\bar{\eta}+\sum_{i=1}^{\mathfrak{k}}\eta_{2i-1}\Upsilon_{2\mathfrak{k}-2i+2}(w)$ has purely imaginary entries. On the other hand, $\eta_{k}$ with even $k$ are real-valued, and they arise as multiples of the columns linear in odd functions $\wp_{1,1,2i-1}(w)$. Thus, $\wp_{1,1,2i-1}(w)$ are purely imaginary. Similar considerations in the case of $g=2\mathfrak{k}$ bring to the same conclusion. ∎
Proposition 5.
Let a hyperelliptic curve of genus $g$ have all real branch points. Then with a choice of cycles as on fig. 1 and the standard not normalized holomorphic differentials (3a), the periods $\omega$ along $\mathfrak{a}$-cycles are real and the periods $\omega^{\prime}$ along $\mathfrak{b}$-cycles are purely imaginary.
Proof.
Let a hyperelliptic curve (2) have the form $-w^{2}+\Lambda(z)=0$, where by $\Lambda(z)$ the polynomial in $z$ is denoted, and the coefficient of the leading term $z^{2g+1}$ equals $1$. If all branch points are real, then $\Lambda(z)>0$ when $z\in(e_{2k-1},e_{2k})$, $k=1$, …, $g$. Thus, the periods $\omega_{k}$, computed from the holomorphic differentials (3a) along $\mathfrak{a}_{k}$ cycles, are real. On the other hand, $\Lambda(z)<0$ when $z\in(e_{2k},e_{2k+1})$, $k=1$, …, $g$. And so the periods $\omega^{\prime}_{k}$, computed from the holomorphic differentials (3a) along $\mathfrak{b}_{k}$ cycles, are purely imaginary.
Let a hyperelliptic curve (52) have the form $-y^{2}+H(z)=0$, where the leading term $z^{2g+1}$ of the polynomial $H(z)$ has an arbitrary real coefficient $\mathrm{b}^{2}$. Then we divide the curve equation by the coefficient $\mathrm{b}^{2}$, and obtain $(y/\mathrm{b})^{2}=\Lambda(z)$. In this case, periods $\omega_{k}$, $\omega^{\prime}_{k}$ are computed from the holomorphic differentials (53a), which coincide with (3a) after the mentioned transformation of the curve.
Finally, suppose, that the leading term of a genus $g$ hyperelliptic curve of the form $-y^{2}+\Lambda(z)=0$ is $z^{2g+2}$, and so all $22g+$ 2 branch points are finite. In this case we numerate branch points by indices $i=0$, …, $12g+$ 1, $e_{0}$ serves as the base-point, and the cut $(e_{2k+1},\infty)$ is replaced by the cut $(e_{2k+1},\infty)\cup(\infty,e_{0})$. If the canonical cycles are defined as before, $\omega_{k}$ are real, and $\omega^{\prime}_{k}$ are purely imaginary. ∎

### v4

@@ 5.3. Summary
The uniformization of the spectral curve is given by (41) in an implicit form. On the other hand, it brings explicit expressions (42) for dynamic variables $\beta_{2n-1}$, $\alpha_{2n}$. Coordinates $u_{1}$ and $u_{3}$ of the Jacobian variety of the spectral curve serve, up to a constant multiple $-\mathrm{b}$, as parameters $\mathrm{x}$ and $\mathrm{t}$ of the stationary and evolutionary flows, correspondingly.
@@ 6. Reality conditions
It is known that an $N$-gap hamiltonian system in terms of variables of separation $(z_{i},w_{i})$ splits into $N$ independent systems. Each system with coordinate $z_{i}$ and momenta $w_{i}$ describes a motion of mass $1/2$ in the potential $-H(z_{i})$, and periodic motion is located between a pair of roots of the potential where $H(z)>0$. Thus, finite trajectories contain branch points as turning points, and so the argument of $\wp_{1,1}$ in (48) attains half-periods. In order to guarantee that $\wp_{1,1}$ is real-valued we assume that all branch points are real.
Further, we work with the spectral curve (30) reduced to the canonical form (3).
Proposition 1.
Let a hyperelliptic curve of genus $g$ have all real branch points. Then with a choice of cycles as on fig. 1 and the standard not normalized holomorphic differentials (4a), all entries of the period matrix $\omega$ are real, and all entries of the period matrix $\omega^{\prime}$ are purely imaginary.
Proof.
Let a hyperelliptic curve (3) have the form $-y^{2}+\Lambda(x)=0$, where by $\Lambda(x)$ the polynomial in $x$ is denoted. If all branch points are real, then $\Lambda(x)>0$ at $x\in(e_{2k-1},e_{2k})$, $k=1$, …, $g$. Thus, the periods $\omega_{k}$, computed from the holomorphic differentials (4a) along $\mathfrak{a}_{k}$ cycles, are real, cf. (59). On the other hand, $\Lambda(x)<0$ at $x\in(e_{2k},e_{2k+1})$, $k=1$, …, $g$. And so the periods $\omega^{\prime}_{k}$, computed from the holomorphic differentials (4a) along $\mathfrak{b}_{k}$ cycles, are purely imaginary, cf. (59).
Let a hyperelliptic curve (39) have the form $-w^{2}+H(z)=0$, where the leading term $z^{2g+1}$ of the polynomial $H(z)$ has an arbitrary real coefficient $\mathrm{b}^{2}$. By the transformation $z\mapsto x$, $w\mapsto\mathrm{b}y$ we find $(w/\mathrm{b})^{2}=\Lambda(z)$. Thus, the holomorphic differentials (40a) produce the same periods $\omega_{k}$, $\omega^{\prime}_{k}$ as (4a).
Finally, suppose, that the leading term of a genus $g$ hyperelliptic curve of the form $-y^{2}+\Lambda(x)=0$ is $x^{2g+2}$, and so all $22g+$ 2 branch points are finite. In this case, we enumerate branch points by indices $i=0$, …, $12g+$ 1, and $e_{0}$ serves as the base-point. The cut $(e_{2g+1},\infty)$ is replaced by $(e_{2g+1},\infty)\cup\{\infty\}\cup(\infty,e_{0})$. If the canonical cycles are defined as on fig. 1, then $\omega_{k}$ are real, and $\omega^{\prime}_{k}$ are purely imaginary. ∎
@@ 6.1. Singularities of Weierstrass-$\wp$-functions
(49) $\begin{split}&\{2k-1\}\sim\tfrac{1}{2}\omega^{\prime}_{k}+\sum_{i=1}^{k-1}\tfrac{1}{2}\omega_{i},\quad\{2k\}\sim\tfrac{1}{2}\omega^{\prime}_{k}+\sum_{i=1}^{k}\tfrac{1}{2}\omega_{i},\quad k=1,\dots,g,\\ &\{2g+1\}\sim\sum_{i=1}^{g}\tfrac{1}{2}\omega_{i}.\end{split}$
Every half-period has the form $\Omega(I)+\Omega^{\prime}(I^{\prime})$, where $\Omega(I)$ is generated from real half-periods $\tfrac{1}{2}\omega_{k}$, $k=1$, …, $g$, and $\Omega^{\prime}(I^{\prime})$ is generated from purely imaginary half-periods $\tfrac{1}{2}\omega^{\prime}_{k}$, $k=1$, …, $g$. Namely,
@@
The sigma function does not vanish at these half-periods. Within the fundamental domain, the only zero of the sigma function is located at $u=0$.
If $g>1$, then each $\mathcal{I}_{0}$ is obtained from $g$ sets $\{\iota\}$ with all indices different. The corresponding $I^{\prime}$ is obtained by taking the union of $g$ subsets from the collection $\mathfrak{I}^{\prime}_{1}$, and dropping indices which occur even number of times. If at least two subsets in this union coincide, then the resulting $I^{\prime}$ is a union of $g-2$ or less number of subsets from $\mathfrak{I}^{\prime}_{1}$. That means, that the subspace $\mathfrak{U}_{\Re}(I^{\prime})$ contains zeros of the sigma function. Thus, the required $\hat{I}^{\prime}$ is obtained by the union of $g$ different and not empty subsets from the collection $\mathfrak{I}^{\prime}_{1}$. This implies $\hat{I}^{\prime}=\{1,2,\dots,g\}$. ∎
@@
$\Omega(\hat{I})=\left\{\begin{array}[]{ll}\sum_{k=1}^{\mathfrak{k}}\tfrac{1}{2}\omega_{2k},&g=2\mathfrak{k}\\ \sum_{k=0}^{\mathfrak{k}}\tfrac{1}{2}\omega_{2k-1}&g=2\mathfrak{k}-1,\end{array}\right.$
@@ 6.2. Hyperelliptic addition law
Let $\Upsilon_{n}(u)$, $n=2$, …, $g+2$, be $g$-component vector-functions of $u\in\operatorname{\mathfrak{J}}$. We introduce a matrix-function $\textsf{Q}(u)=\big{(}\Upsilon_{2}(u)$, $\Upsilon_{3}(u)$, …, $\Upsilon_{g+1}(u)\big{)}$ with entries $(\textsf{q}_{i,j}(u))$, and a vector-function $\bm{q}(u)=\Upsilon_{g+2}(u)$. Actually, on a curve of the form (3)
(51a) $\Upsilon_{2}(u)\equiv\big{(}\textsf{q}_{i,1}(u)\big{)}=\big{(}\wp_{1,2i-1}(u)\big{)},\quad i=1,\dots,g,$
(51b) $\Upsilon_{3}(u)\equiv\big{(}\textsf{q}_{i,2}(u)\big{)}=-\tfrac{1}{2}\big{(}\wp_{1,1,2i-1}(u)\big{)},$
(51c) $\Upsilon_{k+1}(u)\equiv(\textsf{q}_{i,k}(u))=\textsf{q}_{1,k-2}(u)\Upsilon_{2}(u)+\Upsilon^{\circ}_{k-1}(u),\quad k=3,\dots,g+2,$
(51d) where $\text{where }\Upsilon^{\circ}_{k-1}(u)=\big{(}\textsf{q}_{2,k-2}(u),\textsf{q}_{3,k-2}(u),\dots,\textsf{q}_{g,k-2}(u),0\big{)}^{t}.$
Let $\bar{\nu}=(\nu_{g+2},\nu_{g+4},\dots,\nu_{3g})^{t}$, and $\nu=(\nu_{g},\dots,\nu_{2},\nu_{1})^{t}$. Let $u_{\text{I}}$, $u_{\text{II}}$, $u_{\text{III}}\in\operatorname{\mathfrak{J}}$, subject to $u_{\text{I}}+u_{\text{II}}+u_{\text{III}}=0$. Then the system
$\begin{pmatrix}1_{g}&\textsf{Q}(u_{\text{I}})\\ 1_{g}&\textsf{Q}(u_{\text{II}})\\ 1_{g}&\textsf{Q}(u_{\text{III}})\end{pmatrix}\begin{pmatrix}\bar{\nu}\\ \nu\end{pmatrix}=-\begin{pmatrix}\bm{q}(u_{\text{I}})\\ \bm{q}(u_{\text{II}})\\ \bm{q}(u_{\text{III}})\end{pmatrix}$
@@
$\textsf{P}=\big{(}\Upsilon_{2}(u_{\text{I}})+\Upsilon_{2}(u_{\text{II}}),\Upsilon_{4}(u_{\text{I}})+\Upsilon_{4}(u_{\text{II}}),\dots,\Upsilon_{2g}(u_{\text{I}})+\Upsilon_{2g}(u_{\text{II}})\big{)},$
where $\Upsilon_{k}$, $k>g+2$, are computed by the rule (51c). Let a vector $\Pi=(\Pi_{0}$, $\Pi_{1}$, …, $\Pi_{g})$ be defined as follows
@@
and a vector $\mathrm{N}=(\mathrm{N}_{0}$, $\mathrm{N}_{1}$, …, $\mathrm{N}_{g})$, $\mathrm{N}_{0}=-1$, depending on the parity of $g$, be
(52b) $\begin{split}&g=2\mathfrak{k}-1:\qquad\mathrm{N}_{k}=\sum_{i=1}^{k}\nu_{2k-2i+1}\nu_{2i-1}-\sum_{j=0}^{k}\lambda_{2k-2j}\sum_{i=0}^{j}\nu_{2j-2i}\nu_{2i},\\ &g=2\mathfrak{k}:\qquad\quad\ \ \mathrm{N}_{k}=-\sum_{i=0}^{k}\nu_{2k-2i}\nu_{2i}+\sum_{j=1}^{k}\lambda_{2k-2j}\sum_{i=1}^{j}\nu_{2j-2i+1}\nu_{2i-1},\end{split}$
where $\nu_{0}=1$, and $\lambda_{0}=1$. Then addition formulas for $\wp_{1,2i-1}$ are given by
(53) $\wp_{1,2i-1}(u_{\text{III}})=\sum_{j=0}^{i}\mathrm{N}_{i-j}\Pi_{j}.$
They work in an arbitrary genus $g$, and $\nu_{k}$ is supposed to be zero if there is no such entry of $\bar{\nu}$, $\nu$ in this genus. In particular,
(54) $\begin{split}&g=2\mathfrak{k}-1:\qquad\wp_{1,1}(u_{\text{III}})=\nu_{1}^{2}-2\nu_{2}-\lambda_{2}-\wp_{1,1}(u_{\text{I}})-\wp_{1,1}(u_{\text{II}}),\\ &g=2\mathfrak{k}:\qquad\quad\ \ \wp_{1,1}(u_{\text{III}})=\nu_{1}^{2}-2\nu_{2}-\wp_{1,1}(u_{\text{I}})-\wp_{1,1}(u_{\text{II}}),\end{split}$
where $\nu_{2}=0$ in the case of genus $1$.
Expressions for $\nu_{k}$ in terms of $\wp_{1,2i-1}(u_{\text{I}})$, $\wp_{1,2i-1}(u_{\text{II}})$, $\wp_{1,1,2i-1}(u_{\text{I}})$, $\wp_{1,1,2i-1}(u_{\text{II}})$ are obtained from the system
(55) $\begin{pmatrix}1_{g}&\textsf{Q}(u_{\text{I}})\\ 1_{g}&\textsf{Q}(u_{\text{II}})\end{pmatrix}\begin{pmatrix}\bar{\nu}\\ \nu\end{pmatrix}=-\begin{pmatrix}\bm{q}(u_{\text{I}})\\ \bm{q}(u_{\text{II}})\end{pmatrix}.$
Expressions for $\wp_{1,1,2i-1}(u_{\text{III}})$ in terms of $\wp_{1,2i-1}(u_{\text{III}})$ and $\nu_{k}$ are obtained from
(56) $\bar{\nu}+\textsf{Q}(u_{\text{III}})\nu=-\bm{q}(u_{\text{III}}),$
@@
Remark 5.
The addition law is obtained from the entire rational function $\mathcal{R}_{3g}(x,y)=y\nu_{y}(x)+\nu_{x}(x)$ of weight $3g$, where $\nu_{y}(x)=\sum_{i=1}^{[(g-1)/2]}\nu_{g-1-2i}x^{i}$, and $\nu_{x}(x)=\sum_{i=0}^{[3g/2]}\nu_{3g-2i}x^{i}$. With the help of the solution (9) of the Jacobi inversion problem, $y$ is eliminated from $\mathcal{R}_{3g}$, and the degree of $x$ is reduced to $g-1$. Pre-images of $u_{\text{I}}$, $u_{\text{II}}$, $u_{\text{III}}$ are supposed to be zeros of $\mathcal{R}_{3g}$. Thus, coefficients of $\mathcal{R}_{3g}$ reduced to a degree $g-1$ polynomial in $x$ produce $3g$ equations (55), (56). The expressions (53) are derived from the equality
$\nu_{y}^{2}f(x,-\nu_{x}/\nu_{y})=\mathcal{R}_{2g}(x,y;u_{\text{I}})\mathcal{R}_{2g}(x,y;u_{\text{II}})\mathcal{R}_{2g}(x,y;u_{\text{III}}),$
which reflects the fact that $u_{\text{I}}$, $u_{\text{II}}$, $u_{\text{III}}$ form the divisor of zeros of $\mathcal{R}_{3g}$.
@@ 6.3. Real-valued Weierstrass-$\wp$-functions
Through the help of the addition law one can obtain expressions for $\wp_{1,2i-1}(u+\imath\upsilon)$ and $\wp_{1,1,2i-1}(u+\imath\upsilon)$ in terms of $\wp_{1,2i-1}(u)$, $\wp_{1,2i-1}(\imath\upsilon)$, $\wp_{1,1,2i-1}(u)$, $\wp_{1,1,2i-1}(\imath\upsilon)$, $i=1$, …, $g$. The mentioned $2g$ Weierstrass-$\wp$-functions serve as generators in the differential field of all multiply periodic functions on a hyperelliptic curve of genus $g$, see [12]. These $2g$ functions arise in the solution (9) of the Jacobi inversion problem.
Below, we prove that $\wp_{i,j}$ on subspaces $\mathfrak{U}_{\Re}(I^{\prime})$, and $\mathfrak{U}_{\Im}(I)$ are real-valued.
@@
Let $\upsilon\in\operatorname{\mathbb{R}}^{g}$ be fixed, and $\wp_{1,1,2i-1}(\imath\upsilon)=0$, that is $\imath\upsilon$ is a half-period. Then for all $u\in\operatorname{\mathbb{R}}^{g}$, functions $\wp_{1,2i-1}(u+\imath\upsilon)$ and $\wp_{1,1,2i-1}(u+\imath\upsilon)$, $i=1$, …, $g$, are real-valued.
@@
The fact that $\wp_{1,1,2i-1}$ vanishes at all half-periods on a hyperelliptic curve follows immediately from (9b).
Next, recall that $\wp_{i,j}$ are even functions, and so $\wp_{i,j}(\imath\upsilon)=\widehat{\wp}_{i,j}(\upsilon)$ are real-valued. On the other hand, $\wp_{i,j,k}$ are odd, and so $\wp_{i,j,k}(\imath\upsilon)=\imath\widehat{\wp}_{i,j,k}(\upsilon)$ are purely imaginary. Since $\imath$ emerges only from $\wp_{1,1,2i-1}(\imath\upsilon)$, which vanishes according to the condition, all functions $\wp_{1,2i-1}(u+\imath\upsilon)$ and $\wp_{1,1,2i-1}(u+\imath\upsilon)$ are real-valued. ∎
@@
Let $u\in\operatorname{\mathbb{R}}^{g}$ be fixed, and $\wp_{1,1,2i-1}(u)=0$, that is $u$ is a half-period. Then for all $\upsilon\in\operatorname{\mathbb{R}}^{g}$, functions $\wp_{1,2i-1}(u+\imath\upsilon)$, $i=1$, …, $g$, are real-valued, and $\wp_{1,1,2i-1}(u+\imath\upsilon)$, $i=1$, …, $g$, are purely imaginary.
@@
From the structure of Q and $\bm{q}$ in (55) we see the following.
@@
If $g=2\mathfrak{k}-1$, then Q contains $\mathfrak{k}-1$ columns linear in odd functions $\wp_{1,1,2i-1}$, and $\bm{q}$ is linear in $\wp_{1,1,2i-1}$. These columns of $\textsf{Q}(\imath\upsilon)$ and $\bm{q}(\imath\upsilon)$ have purely imaginary entries. Due to $\wp_{1,1,2i-1}(u)=0$, the corresponding columns of $\textsf{Q}(u)$ vanish, as well as $\bm{q}(u)$. The system (55) with $u_{\text{I}}=u$ and $u_{\text{II}}=\imath\upsilon$ is solved by Cramer’s rule. Thus, all entries of $\bar{\nu}$, which are indexed by odd numbers, have purely imaginary values, as well as the entries $\nu_{k}$ of $\nu$ with odd $k$. The entries $\nu_{k}$ with even $k$ are real-valued.
@@
If $g=2\mathfrak{k}$, then Q contains $\mathfrak{k}$ columns linear in odd functions $\wp_{1,1,2i-1}$, and $\bm{q}$ is expressed in terms of even functions $\wp_{1,2i-1}$ only. Thus, all entries of $\bar{\nu}$, which are indexed by even numbers, are real-valued, as well as the entries $\nu_{k}$ of $\nu$ with even $k$. The entries $\nu_{k}$ with odd $k$ are purely imaginary.
Therefore, $\nu_{k}$ with odd $k$ are purely imaginary, and with even $k$ are real-valued, in any genus.
@@
Next, we look at the expressions for $\textsf{N}_{k}$ given by (52b). Evidently, all $\textsf{N}_{k}$ are real-valued, since $\nu_{k}$ with odd $k$ appear in quadratic terms $\nu_{2j-2i+1}\nu_{2i-1}$, which are real-valued. Therefore, expressions for $\wp_{1,2i-1}(u+\imath\upsilon)$ given by (53) with $u+\imath\upsilon=-u_{\text{III}}$ are real-valued.
@@
Expressions for $\wp_{1,1,2i-1}(u+\imath\upsilon)$ are obtained from (56) with $u+\imath\upsilon=-u_{\text{III}}$. If $g=2\mathfrak{k}-1$, then $\nu_{k}$ with odd $k$ are multiplied by real-valued columns expressed in terms of even functions $\wp_{1,2i-1}(u_{\text{III}})$ only. Since $\nu_{k}$ with odd $k$ are purely imaginary, the expression $\bar{\nu}+\sum_{i=1}^{\mathfrak{k}}\nu_{2i-1}\Upsilon_{2\mathfrak{k}-2i+2}(u_{\text{III}})$ has purely imaginary entries. On the other hand, $\nu_{k}$ with even $k$ are real-valued, and they arise as multiples of the columns linear in odd functions $\wp_{1,1,2i-1}(u_{\text{III}})$. Thus, $\wp_{1,1,2i-1}(u_{\text{III}})$ are purely imaginary. Similar considerations in the case of $g=2\mathfrak{k}$ bring to the same conclusion. ∎